Grade 9Magical Mathematics[Interesting Approach]

f(x).f(y)=f(x+y)
find f(x)

Leo Pixie
12 Years agoGrade 9

2 Answers

Arun Kumar
12 Years ago
Hello Student,

\\f(x)f(y)=f(x+y) \\=>f'(x)f(y)=f'(x+y)---eq1 \\y=0 \\=>f(0)=1 \\=>f(x)f(-x)=1----eq2 \\$from 1 & 2$ \\f(x)=e^x

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
yaswanth
11 Years ago
f(x).f(y)=f(x+y)
by trial n error
we can say that
f(x)=ax
it satisfies given condition